OpenStudy (anonymous):

Graph the function 2 3sqrtx+2 +4 (So confused :/ )

OpenStudy (some.random.cool.kid):

do they give you a graph to work with?

OpenStudy (some.random.cool.kid):

or no

OpenStudy (solomonzelman):

You can use, https://www.desmos.com/calculator or if you want, we can discuss the asymptotes, the intercepts, and the restrictions for domain and range (if any)... then graph them by hand, labeling the whatever information we find. (that would be more like it, a school graphing assignment)

OpenStudy (anonymous):

\[2 \sqrt[3]{x+2}+4 \] Better visual and they give me graphs for answer choices

OpenStudy (solomonzelman):

\(\Large\color{ black }{\large {\bbox[5pt, lightgold ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline~~~~~~~~~~~~~~~~~~~~~~~~~~~\textbf{Shifts}~~~~~~~~~~~~~~~~~~~~~~~~~~~&~\bf{c~~~units~~~~} \\ \hline \\f(x)= \sqrt[3]{x} ~~~ ⇒ ~~~ f(x)= \sqrt[3]{x \normalsize\color{blue}{ -~\rm{c}} } &~\rm{to~~the~~right~} \\ \text{ } \\ f(x)= \sqrt[3]{x} ~~~ ⇒ ~~~ f(x)= \sqrt[3]{x \normalsize\color{blue}{ +~\rm{c}} } &~\rm{to~~the~~left ~} \\ \text{ } \\ f(x)= \sqrt[3]{x} ~~~ ⇒ ~~~ f(x)= \sqrt[3]{x} \normalsize\color{blue}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ f(x)= \sqrt[3]{x} ~~~ ⇒ ~~~ f(x)= \sqrt[3]{x} \normalsize\color{blue}{ -~\rm{c} } &~\rm{down~} \\ \\ \hline \end{array} }}}\) And when you say , \(\large\color{black}{ c\sqrt[3]{x} }\) you are stretching the graph of \(\large\color{black}{ \sqrt[3]{x} }\), by a scale factor of c.

OpenStudy (solomonzelman):

opps you have to refresh, sorry.

OpenStudy (some.random.cool.kid):

solo

OpenStudy (solomonzelman):

what?

OpenStudy (some.random.cool.kid):

i need help

OpenStudy (some.random.cool.kid):

on mine

OpenStudy (solomonzelman):

You can tag me, but please not in this question....

OpenStudy (some.random.cool.kid):

i did tag you

OpenStudy (some.random.cool.kid):

you didnt respond so i came back

OpenStudy (solomonzelman):

I can't respond to every single tag right away, you know ? :)

OpenStudy (some.random.cool.kid):

lol :b